Twenty pounds of a salt-water solution contains fifteen percent salt. How much water must be added to weaken the solution to twelve percent salt?
Answer by oscargut(2103)

Salt= 20(0,15)=3 pounds, you must add x pounds of water where:
3/(20+x)=0,12 then 20+x=3/(0,12)=25 then x=5
You must add 5 pounds of water